# Multi-Robot Rendezvous in Unknown Environment with Limited Communication

### v1

@@ III-C Partition of the Space
Multiple estimations for RP can be obtained during PIER. Each estimation can form an edge in $\mathcal{G}_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$. Initially, $\mathcal{E}_{r}$ is an empty set. When single estimations are gradually added to $\mathcal{E}_{r}$, $\mathcal{G}_{r}$ will become partially connected. Assuming that $\mathcal{G}_{r}$ has $m$ connected components, which can be represented by $\mathcal{X}_{i},i=1,2\cdots m$. Based on the definition of connected components, we can be obtained that
@@
For any two robots in the same connected component, their RP is known. If they are not in the same connected component, their RP is unknown. To obtain the RPs in $\mathcal{X}_{i}$, the Pose Graph Optimization (PGO) can be performed using
@@
Collaborative exploration can be performed based on the estimated RPs. For any connected component $\mathcal{X}_{i}$, partition of the space is executed among robots in $\mathcal{X}_{i}$. Similarly with [34], Voronoi diagram is used for the partition of space. The positions of robots in $\mathcal{X}_{i}$ are set as base points for this Voronoi diagram. If $\mathcal{R}^{k}\in\mathcal{X}_{i}$, then $S^{k}$ for $\mathcal{R}^{k}$ can be obtained by
@@
where $\mathbf{p}$ is a point in the Voronoi diagram partition for $\mathcal{R}^{k}$. $\hat{\mathbf{p}}^{k}$ is the position of $\mathcal{R}^{k}$ in the reference frame. $D(\cdot)$ denotes the distance function between $\mathbf{p}$ and robot position. In this paper, Euclidean distance $D(\mathbf{p},\hat{\mathbf{p}}^{k})=||\mathbf{p}-\hat{\mathbf{p}}^{k}||_{2}$ is used. The partition of the space also needs to be dynamically adjusted during the exploration. We set the criteria for re-partitioning as follows: when new estimations for RP are obtained or the region assigned to any robot has been fully explored.
@@ III-D Algorithm for PIER
Input: Graph for RP $\mathcal{G}_{r}$, communication graph $\mathcal{G}_{c}$,$n$ robots
1 $\mathcal{E}_{r}=\varnothing$
@@ IV Optimal Rendezvous Point Selection
In the FHT-Map, each node $v_{f}^{k,i}$ stores a convex local free space $s^{k,i}$ around it. $\hat{\textbf{p}}^{*}$ may exist in any of the $s^{k,i}$. Besides, since the problem is non-convex, we solve it by an approach based on sampling. Points are sampled in each $s^{k,i}$ with a density of $\delta$. Then, based on path planning methods on the FHT-Map [24], we can obtain the value $F(\cdot)=\max_{k=1\cdots n}D(\tilde{\textbf{p}}^{k}(t_{0}),\cdot)$. Finally, the minimal $F(\cdot)$ of the sampled points will be selected as $\hat{\textbf{p}}^{*}$. The algorithm for selecting sub-optimal rendezvous point $\hat{\textbf{p}}^{*}$ is presented in Algorithm 3 based on the merged topological map.
@@ 6       $L\leftarrow L\cup L^{\prime}$
The main time-consuming step in Algorithm 3 is CalculateF() (Line 1 and 5). Firstly, path length from $\tilde{\textbf{p}}^{k}(t_{0})$ to each node $\{\textbf{p}^{k,i}\}$ (Line 1) can be quickly obtained in the FHT-Map. Besides, path planning for points $\mathbf{p}^{\text{PointSet}}$ in $s^{k,i}$ (Line 5) consists of path on FHT-Map and path from nodes to $\mathbf{p}^{\text{PointSet}}$, where any actual path planning algorithm is not required. So, our proposed method is very time-efficient.
@@ V-A1 Simulation Setup
Simulations are conducted in the ROS-Gazebo simulator with AMD Ryzen 3900X CPU at 4.00 GHz (24 cores), 64 GB RAM, and NVIDIA RTX 2080 Ti GPU. Turtlebot Burger robots are used in this work, and each robot is equipped with an RPLIDAR A2 LIDAR (sensing range 12 m), which is the same for real-world experiments. Cartographer is used for localization during exploration. The robot uses four RGB cameras which are arranged horizontally in a circular manner without overlap of Field of View (FOV) to capture information from all directions. Each camera has a FOV of $90^{\circ}$, with a resolution of 640$\times$480. Two simulation environments are used in this work, where small is 485 $\text{m}^{2}$ and office is 3285.22 $\text{m}^{2}$. In all simulations, $\gamma_{1}$ and $\gamma_{2}$ are set to $\exp(-0.5)$ and $\exp(-1)$, $\omega_{1}$ and $\omega_{2}$ are set to 1 and 2, $th_{\text{sim}}$ is set to 0.94, $th_{\text{match}}$ is 0.96, and $\sigma_{c}$ is set to 4.
@@ V-A2 Real-world Experiment
Real-world experiment is conducted using Turtlebot Burger robots equipped with a NVIDIA Jetson TX2 and a Realsense D435i camera. Besides, we set $\sigma_{c}$ to 1, while keeping other hyper-parameters constant.
@@ V-C Comparison and Ablation Studies
To validate the proposed method, comparison and ablation studies are conducted. Firstly, we compare our overall framework with two methods. In GM, an RRT-based [38] strategy is used for exploration and grid maps are transmitted for RP estimation. Besides, 10 points are randomly sampled in the merged map and the point with minimal $F(\cdot)$ is selected for rendezvous. In TM, the robot exploration and RP estimation is based on traditional topological maps [23], and the rendezvous point is selected among all nodes. The time when exploration is finished is defined as $\mathbf{t_{0}}$, the time when rendezvous is achieved is defined as $\mathbf{t_{1}}$, and the volume of transmitted data is also presented in Table II.
Compared with GM, we reduce the exploration time $\mathbf{t_{0}}$ by 29.15 % and 62.11 % in small and office respectively. Besides, due to only a lightweight topological map is transmitted in this work, we reduce the communication burden by 91.50 % and 96.02 %. Compared with TM, we reduce the time for achieving rendezvous $\mathbf{t_{1}}$ by 46.43 % and 33.99 %, which can be attributed to that our method accelerates the acquisition of RP and allows for the selection of better rendezvous points.
Ablation studies are conducted to validate the effectiveness of the exploration strategy and the proposed main node selection module. As for exploration strategy, we compare PIER against RRT-based [38] (RRT) and m-TSP-based [33] (TSP) exploration. Detailed results are presented in Table II. PIER outperforms the other two methods in terms of exploration time and transmitted data. Compared with RRT and TSP, PIER reduces the $\mathbf{t_{0}}$ by 30.32 % and 49.08 % in small (29.44 % and 41.20 % in office). Due to exploration time is reduced, a smaller map is constructed, which also results in a decrease in the amount of transmitted data. Besides, we remove heat value evaluation in main nodes selection (NS), and only $\mathcal{D}(\textbf{p}^{k}(\tau))$ is considered in Eq. 2. The introduction of heat value evaluation can accelerate the exploration by 13.47% and 16.83% in two environments.
TABLE II: Comparison with different exploration methods. Methods with “*” and “#” are comparison and ablation experiments respectively.
@@ V-D Evaluation on Rendezvous Point Selection
Figure 4: Illustration of the selected rendezvous points. The color in the map represents $F(\textbf{p})$ at location p obtained by A* algorithm. The red circles represent the positions of the robots. The yellow star indicates the locations of the global optimal points $\textbf{p}^{*}$. The plus sign represents the position of $\hat{\textbf{p}}^{*}$ based on FHT-Maps. (a) Two robots. (b) Four robots. (c) Six robots. (d) Eight robots.
We also compare the efficiency of rendezvous point computation based on the topological map and the grid map. For our method, the computation time is influenced by the sampling density $\delta$. For the method based on the grid map, we randomly sample $m$ points in free space and use the A* algorithm to obtain $F(\cdot)$, and the point with minimum $F(\cdot)$ is selected as the rendezvous point. The main factor affecting the computation time is $m$.
Fig. 5 reports the relationship between computation time and error $F(\hat{\textbf{p}}^{*})-F(\textbf{p}^{*})$. For the FHT-Map with support node density of 2 m (red line), compared to the grid map, it achieves a result closer to the optimal solution at a low computation time. However, as $\delta$ decreases further, it cannot obtain a more accurate result. It is because the topological map is a sparse representation of the environment, which prevents obtaining accurate shortest distances between two points. However, by using an FHT-Map with denser nodes (node density is 1 m, brown line), we are able to obtain a result that is closer to the theoretically optimal rendezvous point.
Figure 5: Comparison on time consuming and error based on different map. A near optimal solution $\hat{\textbf{p}}^{*}$ can be obtained on FHT-Map much faster than on grid map. Besides, if there are denser nodes in FHT-Map, a more accurate result can be obtained.
V-E Real World Experiment
To validate the effectiveness of our method, real-world experiment is conducted in an indoor environment with size 5.5 $\times$ 9.3  m2 using two robots. A snapshot of the result is shown in Fig. 6. It takes 91 s for PIER and 57 s to achieve rendezvous in this case. The total amount of transmitted data is 26.71 KB, which indicates that rendezvous is achieved with limited communication.
@@ VI Conclusion & Future Work
In this work, we present a solution for the rendezvous problem in unknown environments with limited communication. Firstly, PIER is proposed for the task of incomplete exploration. Then, each robot constructs and shares an FHT-Map during exploration, and RP estimation can be realized using FHT-Maps. Finally, when $\mathcal{G}_{r}$ becomes a connected graph, a sub-optimal rendezvous point will be selected. Simulations are conducted to validate the effectiveness of the proposed method, where a shorter time for exploration and calculation of $\hat{\textbf{p}}^{*}$ is achieved.

### v2

@@ III-C Partition of the Space
Multiple estimations for RP can be obtained during PIER. Any edge in $\mathcal{G}_{r}$ indicates that there is at least one estimated RPs. Initially, $\mathcal{E}_{r}$ is an empty set. When single estimations are gradually added to $\mathcal{E}_{r}$, $\mathcal{G}_{r}$ will become partially connected. Assuming that $\mathcal{G}_{r}$ has $m$ connected components, which can be represented by $\mathcal{X}_{i},i=1,2\cdots m$. Based on the definition of connected components, we can be obtained that
@@
For any two robots in the same connected component, their RP is known. If they are not in the same connected component, their RP is unknown. To obtain the RPs in $\mathcal{X}_{i}$, the PGO can be performed using
@@
Collaborative exploration can be performed based on the estimated RPs. For any connected component $\mathcal{X}_{i}$, partition of the space is executed among robots in $\mathcal{X}_{i}$. Similarly with [29], Voronoi diagram is used for the partition of space and the collaborative behavior of robots is enhanced. The positions of robots in $\mathcal{X}_{i}$ are set as base points for this Voronoi diagram. If $\mathcal{R}^{k}\in\mathcal{X}_{i}$, then $S^{k}$ for $\mathcal{R}^{k}$ can be obtained by
@@
where $\mathbf{p}$ is a point in the Voronoi diagram partition for $\mathcal{R}^{k}$. $\hat{\mathbf{p}}^{k}$ is the position of $\mathcal{R}^{k}$ in the reference frame. $D(\cdot)$ denotes the distance function between $\mathbf{p}$ and robot position. In this paper, Euclidean distance $D(\mathbf{p},\hat{\mathbf{p}}^{k})=||\mathbf{p}-\hat{\mathbf{p}}^{k}||_{2}$ is used. The partition of the space also needs to be dynamically adjusted during the exploration. We set the criteria for re-partitioning as follows: when new estimations for RP are obtained or the region assigned to any robot has been fully explored. Besides, considering that each robot has same global information using $\mathcal{G}_{c}$, for example, estimated RPs, each robot can perform PGO to obtain the refined poses of other robots. Then, each robot can execute the process of space partition independently and decide where to explore next.
@@ III-D Algorithm for PIER
Input: Graph for RP $\mathcal{G}_{r}$, communication graph $\mathcal{G}_{c}$
@@ IV Optimal Rendezvous Point Selection
In the FHT-Map, each node $v_{f}^{k,i}$ stores a convex local free space $s^{k,i}$. $\hat{\textbf{p}}^{*}$ may exist in any of the $s^{k,i}$. Besides, since the problem is non-convex, we solve it by an approach based on sampling. Points can be sampled in each $s^{k,i}$ with a density of $\delta$. Then, based on path planning methods on the FHT-Map [21], we can obtain the value $F(\cdot)=\max_{k=1\cdots n}D(\tilde{\textbf{p}}^{k}(t_{0}),\cdot)$ for each point. Finally, the minimal $F(\cdot)$ of sampled points will be selected as $\hat{\textbf{p}}^{*}$. The algorithm for selecting sub-optimal rendezvous point $\hat{\textbf{p}}^{*}$ is presented in Algorithm 3 based on the merged topological map.
@@ 6       $L\leftarrow L\cup L^{\prime}$
Due to that CalculateF() (Line 1 and 5), which means calculate $F(\cdot)$ for given points, involves virtual path planning on the topological maps, it is the main time-consuming step in Algorithm 3. However, the time spent on CalculateF() can be reduced. Firstly, path length from $\tilde{\textbf{p}}^{k}(t_{0})$ to each node $\{\textbf{p}^{k,i}\}$ (Line 1) can be quickly obtained in the FHT-Map. Besides, path planning for points $\mathbf{p}^{\text{PointSet}}$ in $s^{k,i}$ (Line 5) consists of path on FHT-Map and path from nodes to $\mathbf{p}^{\text{PointSet}}$, where any actual path planning algorithm is not required. So, our proposed method is very time-efficient.
@@ V-A1 Simulation Setup
Simulations are conducted in the ROS-Gazebo simulator with AMD Ryzen 3900X CPU at 4.00 GHz (24 cores), 64 GB RAM, and NVIDIA RTX 2080 Ti GPU. Turtlebot Burger robots are used in this work, and each robot is equipped with an RPLIDAR A2 LIDAR (sensing range 12 m), which is the same for real-world experiments. Cartographer is used for localization during exploration. The robot uses four RGB cameras which are arranged horizontally in a circular manner without overlap of Field of View (FOV) to capture information from all directions. Each camera has a FOV of $90^{\circ}$, with a resolution of 640$\times$480. Two simulation environments are used in this work, where small is 485 $\text{m}^{2}$ and office is 3285.22 $\text{m}^{2}$. In all simulations, $\gamma_{1}$ and $\gamma_{2}$ are set to $\exp(-0.5)$ and $\exp(-1)$, $\omega_{1}$ and $\omega_{2}$ are set to 1 and 2, $\theta_{\text{sim}}$ is set to 0.94, $\theta_{\text{match}}$ is 0.96, and $\sigma_{c}$ is set to 4.
@@ V-A2 Real-world Experiment
To validate the effectiveness of our method, real-world experiments are conducted in an indoor environment with size 5.5 $\times$ 9.3  m2 using two robots. Turtlebot Burger robots equipped with a NVIDIA Jetson TX2 and a Realsense D435i camera are used in the experiments. Besides, due to the size of the real-world scene is small, to guarantee that enough number of main nodes can be established for RP estimation, we set $\sigma_{c}$ to 1, while keeping other hyper-parameters constant.
@@ V-C Comparison and Ablation Studies
To validate the proposed method, comparison and ablation studies are conducted. We compare our overall framework with two methods. The first one transfers grid maps (GM) between robots and the second one transfers traditional topological maps (TM). In GM, the next goal for $\mathcal{R}^{k}$ is selected as the frontier among all frontiers in $\mathcal{X}(\mathcal{R}^{k})$ with maximum utility based on [33]. RP estimation is achieved based on ROS package Multirobot_Map_Merge. Besides, 10 points are randomly sampled in the merged map and the point with minimal $F(\cdot)$ is selected for rendezvous. In TM, the robot exploration and RP estimation is based on traditional topological maps [20], and the rendezvous point is selected among all nodes. In these methods, once the RPs between all robots are determined, rendezvous will be performed. The time when exploration is finished is defined as $\mathbf{t_{0}}$, the time when rendezvous is achieved is defined as $\mathbf{t_{1}}$, and the volume of transmitted data is also presented in Table II.
Compared with GM, we reduce $\mathbf{t_{0}}$ by 29.15 % and 62.11 % in small and office respectively. Besides, due to only a lightweight topological map is transmitted in this work, we reduce the communication burden by 91.50 % and 96.02 %. Compared with TM, we reduce the time for achieving rendezvous $\mathbf{t_{1}}$ by 46.43 % and 33.99 %. It can be attributed to that PIER and transferred $\boldsymbol{\psi}^{k,i}$ can accelerate the acquisition of RP. And the more detailed representation of the environmental structure allows for a better selection rendezvous point. Besides, the average data upload per robot per second is 0.493 KB based on FHT-Map in office. With the bandwidth of 1 Mbps, the latency introduced is 3.943 milliseconds, which has minor influence on the system.
Ablation studies are conducted to validate the effectiveness of the exploration strategy and the proposed main node selection module. As for exploration strategy, we compare PIER against RRT-based [33] (RRT), which is the same as GM, and m-TSP-based [28] (TSP) exploration. Detailed results are presented in Table II. PIER outperforms the other two methods in terms of exploration time and transmitted data. Compared with RRT and TSP, PIER reduces the $\mathbf{t_{0}}$ by 30.32 % and 49.08 % in small (29.44 % and 41.20 % in office). Due to exploration time is reduced, a smaller map is constructed, which also results in a decrease in the amount of transmitted data. Besides, we remove heat value evaluation in main nodes selection (NS), and only $\mathcal{D}(\textbf{p}^{k}(\tau))$ is considered in Eq. 2. The introduction of heat value evaluation can accelerate the exploration by 13.47% and 16.83% in two environments.
TABLE II: Comparison and ablation studies. Methods with “*” and “#” are comparison and ablation experiments respectively.
@@
Three hyper-parameters $\gamma_{1}$, $\gamma_{2}$, and $\sigma_{c}$ are used in the main node selection of experiments. A smaller $\sigma_{c}$ or a larger $\gamma_{2}$ will contribute to an FHT-Map with more main nodes. It can improve the probability of finding matched main nodes, but it leads to an increase in communication burden. $\gamma_{1}$ influences the ratio of potential main nodes.
@@ V-D Evaluation on Rendezvous Point Selection
Figure 4: Illustration of the selected rendezvous points. The color in the map represents $F(\textbf{p})$ at location p obtained by A* algorithm. The red circles represent robots’ positions. The yellow star indicates the global optimal points $\textbf{p}^{*}$. The plus sign represents the position of $\hat{\textbf{p}}^{*}$ based on FHT-Maps. (a) Two robots. (b) Four robots. (c) Six robots. (d) Eight robots.
@@
Figure 5: Comparison on time consumption and error based on different map. A near optimal solution $\hat{\textbf{p}}^{*}$ can be obtained on FHT-Map. Besides, if there are denser nodes in FHT-Map, a more accurate result can be obtained.
We also compare the efficiency of rendezvous point computation based on the topological map and the grid map. For our method, the computation time is influenced by the sampling density $\delta$. For the method based on the grid map, we randomly sample $m$ points in free space and use the A* algorithm to obtain $F(\cdot)$, and the point with minimum $F(\cdot)$ is selected as the rendezvous point. The main factor affecting the computation time is $m$. Fig. 5 reports the relationship between computation time and error $F(\hat{\textbf{p}}^{*})-F(\textbf{p}^{*})$. For the FHT-Map with support node density of 2 m (red line), compared to the grid map, it achieves a result closer to the optimal solution at a low computation time. However, as $\delta$ decreases further, it cannot obtain a more accurate result. It is because the topological map is a sparse representation of the environment, which prevents obtaining accurate shortest distances between two points. However, by using an FHT-Map with denser nodes (node density is 1 m, brown line), we are able to obtain a result that is closer to the theoretically optimal rendezvous point.
V-E Real-world Experiment
A snapshot of the real-world experiment is shown in Fig. 6. It takes 91 s for PIER and 57 s to achieve rendezvous. WiFi is used for transmitting data. The total amount of transmitted data is 26.71 KB, which indicates that rendezvous is achieved within constrained of communication bandwidth.
@@ VI Conclusion & Future Work
In this work, we present a solution for the rendezvous problem in unknown environments with limited communication. Firstly, PIER is proposed for the task of incomplete exploration. Then, each robot constructs and shares an FHT-Map during exploration, and RP estimation can be realized based on this map. Finally, when $\mathcal{G}_{r}$ becomes a connected graph, a sub-optimal rendezvous point will be selected. Simulations are conducted to validate the effectiveness of the proposed method, where a shorter time for exploration and calculation of $\hat{\textbf{p}}^{*}$ is achieved.